The equilibrium constant for the chemical equationI've tried this a million times and I keep getting it wrong!
Answer #1

Kp = Kc* (RT) ^(delta n )

delta n= 2-4 = -2

so,

Kc = Kp* (RT)^2

Kc= ( 4.46)* ( (273+197)*8.314)^2

Kc = 68100549